An air bubble from the bottom of a lake becomes triple itself as it reaches on the water surface. How deep is the lake?

Depth = _____ meters


1
Expert's answer
2021-09-14T10:58:42-0400

We know that

"P_sV_s=P_dV_d"

Assuming atmospheric pressure to be 75 cm of Hg and the density of water to be 1/10 of the density of mercury, the depth of the lake.

"P_s=1atm\\P_d=1+\frac{h}{10}"

"V_s=3V\\V_d=V"

Put value

"1\times 3V=(1+\frac{h}{10}) \times V"

"3=1+\frac{h}{10}"

"2=\frac{h}{10}\\h=20m"

Depth(h)=20m